OpenStudy (anonymous):

If x2 + y2 = 100, and xy = 18, find the value of x - y.

OpenStudy (kamille):

Firstly, you need to solve the system of equations. Do you know how to do it?

OpenStudy (akashdeepdeb):

Or use (x-y)2 = x2 + y2 - 2xy = 100 - 36 = 64 Now (x-y) = rt 64 = 8 !! :D